OpenStudy (anonymous):

Use the point-slope formula to write an equation of the line givn the following information. The line passes thru the points (-2,5) and (0,1)

jigglypuff314 (jigglypuff314):

\[slope = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] so \[\frac{ 5-1 }{ -2-0 }\] which give you your slope then you plug that into the point-slope form which is y-y1=m(x-x1)

OpenStudy (anonymous):

Okay , so I get the slope then for the y1 i'd but the 5 and then in x1 the -2?